# CryoBench: Diverse and challenging datasets for the heterogeneity problem in cryo-EM

### v1

@@ C.4 Additional Qualitative Results
For additional qualitative evaluation, we provide visualizations of the reconstructed volumes and latent spaces for each method and dataset in CryoBench. Figure S12 shows the latent spaces colored by dihedral angle (CV1) and center of mass distance (CV2) for IgG-RL dataset. Figures S13-S19 display volumes from $k$-means cluster centers, with points in the UMAP visualization of the latent space corresponding to each center.
@@ C.5 Additional Information Imbalance Results
CTF and Pose: Information imbalance with respect to the ground truth latent pose (rotation only, not translation) and CTF parameters is generally in the orthogonal region (1,1) for all methods (Figs. S20, S21). However, zooming in, for pose, CryoDRGN and Opus-DSD are off the shared information x=y line, indicating their minor entanglement is more pronounced than other methods. For CTF, the trends are less clear, but Opus-DSD and 3D Class abinit are generally the furthest away from the orthogonal region.
@@ C.6 Additional Spike-MD embedding analysis
Neighborhood Similarity and Information Imbalance plots for Spike-MD are shown in Figure S22. We observe a relatively low similarity in neighborhoods between the embeddings and the ground truth molecular dynamics collective variables for small neighbhoorhood radii, consistent with qualitative observations from UMAP visualizations (Fig. 5). Information imbalance of the Spike-MD dataset (Fig. S22-right) shows 3DVA on the shared information line at (0.5,0.5) - a very similar result as in IgG-1D. Opus-DSD and CryoDRGN2 are near (0.9,0.6), the closest to the orthogonal region for the Spike-MD dataset compared with other methods. For Opus-DSD, this is the closest to the orthogonal region compared with its information imbalance on the other datasets. For CryoDRGN2, this is a similar value as the challenging datasets (IgG-RL and Tomotwin-100). The other methods employed in these experiments (CryoDRGN, DRGN-AI-fixed, 3DFlex, RECOVAR, DRGN-AI) are closer to the equivalent zone and cluster together near (0.5,0.2).
@@ D.2 Heterogeneity
  • 3D Structure: The spatial form or shape of an object, which in the context of cryo-EM refers to the 3D structure of biomolecules. Often contrasted with the sequence of a biomolecule, or schematic (e.g. 2D) representations communicating atom type of bond connectivity.
@@ D.3 Model and Representation
  • 3D Map, Volume, Density, Model: A representation of spatial data, in cryo-EM this typically refers to the 3D Coulombic (electric, electrostatic) potential instead of the electron density in other structural biology techniques based on X-ray diffraction. [65, 66]

### v2

@@ C.4 Additional Qualitative Results
For additional qualitative evaluation, we provide visualizations of the reconstructed volumes and latent spaces for each method and dataset in CryoBench. Figure S14 shows the latent spaces colored by dihedral angle (CV1) and center of mass distance (CV2) for IgG-RL dataset. Figures S15-S21 display volumes from $k$-means cluster centers, with points in the UMAP visualization of the latent space corresponding to each center.
@@ C.5 Additional Information Imbalance Results
CTF and Pose: Information imbalance with respect to the ground truth latent pose (rotation only, not translation) and CTF parameters is generally in the orthogonal region (1,1) for all methods (Figs. S22, S23). However, zooming in, for pose, CryoDRGN and Opus-DSD are off the shared information x=y line, indicating their minor entanglement is more pronounced than other methods. For CTF, the trends are less clear, but Opus-DSD and 3D Class abinit are generally the furthest away from the orthogonal region.
@@ C.6 Additional Spike-MD embedding analysis
Neighborhood Similarity and Information Imbalance plots for Spike-MD are shown in Figure S24. We observe a relatively low similarity in neighborhoods between the embeddings and the ground truth molecular dynamics collective variables for small neighbhoorhood radii, consistent with qualitative observations from UMAP visualizations (Fig. 5). Information imbalance of the Spike-MD dataset (Fig. S24-right) shows 3DVA on the shared information line at (0.5,0.5) - a very similar result as in IgG-1D. Opus-DSD and CryoDRGN2 are near (0.9,0.6), the closest to the orthogonal region for the Spike-MD dataset compared with other methods. For Opus-DSD, this is the closest to the orthogonal region compared with its information imbalance on the other datasets. For CryoDRGN2, this is a similar value as the challenging datasets (IgG-RL and Tomotwin-100). The other methods employed in these experiments (CryoDRGN, CryoDRGN-AI-fixed, 3DFlex, RECOVAR, CryoDRGN-AI) are closer to the equivalent zone and cluster together near (0.5,0.2).
C.7 Pose Error
We compute the pose error for ab initio methods (Table S7). CryoDRGN-AI generally exhibits the lowest pose error. The distribution of pose errors are plotted in Figure S25-S31.
Method IgG-1D IgG-1D noisier IgG-1D noisiest IgG-RL Ribosembly Tomotwin-100 Spike-MD
Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans Rot Trans
CryoDRGN2 6.601 1.056 6.282 1.035 55.876 2.707 6.546 2.527 1.847 0.728 109.581 2.156 3.457 2.022
CryoDRGN-AI 5.216 0.412 5.884 0.435 82.657 4.264 2.712 0.215 2.032 0.872 110.987 5.347 1.562 0.265
3D Class abinit 3.403 0.933 4.312 1.009 82.621 3.583 4.658 0.742 2.834 1.492 113.629 8.980 1.783 0.489
Table S7: Pose Error. Errors are quantified by the median rotation and translation errors compared to the ground truth image poses after global reference frame alignment. The rotation errors are defined as the median of the Geodesic error in units of degrees and the translation errors are defined as the median of the L2 distance in units of pixels after alignment. For the Tomotwin-100 dataset, each structure was aligned separately for all 100 different structures.
Figure S25: Pose error for IgG-1D. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
Figure S26: Pose error for IgG-1D noisier. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
Figure S27: Pose error for IgG-1D noisiest. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
Figure S28: Pose error for IgG-RL. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
Figure S29: Pose error for Spike-MD. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
Figure S30: Pose error for Ribosembly. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
Figure S31: Pose error for Tomotwin-100. Histogram of rotation and translation errors. The first row shows rotation errors, and the second row shows translation errors.
@@ D.3 Model and Representation
  • 3D Map, Volume, Density: A representation of structural data, in cryo-EM this typically refers to the 3D Coulombic (electric, electrostatic) potential instead of the electron density in other structural biology techniques based on X-ray diffraction. [65, 66]